OpenStudy (anonymous):

WHAT IS THE SIMPLIFIED FORM OF EACH EXPRESSION? -(8d - 3w)

OpenStudy (mertsj):

The negative sign in front of the parenthesis means "change everything inside the parenthesis to its opposite."

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

so its -8d-3w ?

OpenStudy (anonymous):

Use Distributive Property: Ex: a(b + c) = ab + ac -(8d - 3w) Is the same as: -1(8d - 3w) So: -1(8d - 3w) = -1(8d) + (-1)(-3w) What's -1 * 8d?

OpenStudy (anonymous):

stillll confused

OpenStudy (anonymous):

;(

OpenStudy (anonymous):